OpenStudy (anonymous):

solve the given problem Use the method of Fourier integrals to find the solution of initial value problem with the partial differential equation. pu/pt=c2p2u/p2t with initial condition u(x,0)=f(x) note here p=partial derivative and 2=square

OpenStudy (anonymous):

So you're looking at: \[\frac{\partial u}{\partial t}-c^2 \frac{\partial^2 u}{\partial t^2}=0\]? Or is the second partial with respect to x?